4.4 A sample with a concentration of 0.3 g/mL was placed in a cell with a length of 5 cm. The resulting rotation at the sodium D line was +1.52°. What is the [α]D?
1
Expert's answer
2019-09-05T07:48:39-0400

The equation for the calculation of "[\alpha]_D^{20}" is:

"[\alpha]_D^{20}=\frac{\alpha \cdot 100}{l \cdot c}"

where "\alpha" is the rotation angle in degrees, l is pathlength (in dm), and c is concentration in grams per 100 mL of solution. Thus,

"[\alpha]_D^{20}=\frac{+1.52\degree \cdot 100}{0.5 dm \cdot 30\frac{g}{100mL}}=+10.13\degree ml \cdot dm^{-1} \cdot g^{-1}"